Calculate the mass in grams of 6.71 x 10^25 molecules of water


1
Expert's answer
2021-03-15T09:29:29-0400

1 mole of water would have "6.02 \u00d710^{23}" (avogadro's constant)


Molar mass of water = 1(2)+16 = 18


Hence


"6.71\u00d710^{25}" molecules of water would have a mass of:

"(\frac{18}{6.02×10^{23}} ) 6.71 ×10^{25} = 2006.3 g"